A card is thrown at random from a well shuffle pack of 52 playing cards. Find the probability that the drawn card is a) neither a king nor a queen b) a number card.
Answer by rothauserc(4718)

a) there are 4 kings and 4 queens in a deck of 52 playing cards
:
Probability neither a king nor a queen elected is (52-8)/52 = 44/52 = 11/13
:
b) there are 4 suits of 9(2-10) number cards = 36 number cards
:
Probability a number card is drawn = 36/52 = 9/13
: